SLRI Machine Status Accelerator Operation Cryogenics
  Operation E-Logbook, Page 428 of 430  Not logged in Accelerator E-Logbook
ID Date Author Event Time Type Section Icon Subject Status RTime (min) Text Attachments
  83   Mon 05 Mar 2012 11:18:53 Natthawut Suradet Mon 05 Mar 2012 11:15:55 RoutineOperation Note Ramping complete   Store beam @ 138 mA
STR-RF-FW Power = 15.4 kW
STR-RF-Ref Power = 216 W
  
  82   Mon 05 Mar 2012 11:17:46 Natthawut Suradet Mon 05 Mar 2012 10:16:38 ConfigurationLINAC Note Adjust Gun Heater   Adjust gun heater 6.15V to 7.56V   
  81   Mon 05 Mar 2012 11:15:43 Natthawut Suradet Mon 05 Mar 2012 10:28:19 Measured ParameterLINAC Note GUN HV pulser     GUN_HV_pulser_(SYN)_5_03_2012.png 
  80   Mon 05 Mar 2012 11:14:01 Natthawut Suradet Mon 05 Mar 2012 11:13:12 RoutineOperation Note Start ramp energy      
  79   Mon 05 Mar 2012 11:12:57 Natthawut Suradet Mon 05 Mar 2012 11:12:29 RoutineOperation Note Stop Injection   @ 140.3 mA   
  78   Mon 05 Mar 2012 11:00:11 Natthawut Suradet Mon 05 Mar 2012 10:50:36 RoutineOperation Note Start Injection  Start injection to STR   
  77   Mon 05 Mar 2012 10:00:59 Sudjakom Jummunt Mon 05 Mar 2012 10:00:37 RepairOperation Note Stop Injection      
  76   Mon 05 Mar 2012 09:24:20 Sudjakom Jummunt Mon 05 Mar 2012 07:30:40 ProblemOther Note Dump beamFixed Water Leak @BL3 
Dump beam @ 83mA
  
  75   Mon 05 Mar 2012 09:06:14 Sudjakom Jummunt Mon 05 Mar 2012 09:04:03 RepairOperation Note Warm up magnets   Warm up HBT,STR magnets   
  74   Mon 05 Mar 2012 03:16:19 Ekwarong Amornwittawat Mon 05 Mar 2012 03:13:19 RoutineBeam Dynamics Note SPS Close Orbit correction   Before - dRMSx     0.1389
     mm
         
 2012-03-05_03-13.PNG 
  73   Mon 05 Mar 2012 03:13:06 Ekwarong Amornwittawat Mon 05 Mar 2012 03:04:33 RoutineOperation Note Ramping energy to 1.2Gev   Ramping complete @ 134mA   
  72   Mon 05 Mar 2012 03:12:30 Ekwarong Amornwittawat Mon 05 Mar 2012 03:02:05 RoutineOperation Note Stop Injection   Stop Injection @135ma   
  71   Mon 05 Mar 2012 00:13:23 Somkiat Wisaiphion Sun 04 Mar 2012 00:00:00 RepairLINAC Note Linac   Still Checking   
  70   Mon 05 Mar 2012 00:04:20 Somkiat Wisaiphion Sun 04 Mar 2012 18:30:00 RepairLINAC Note Linac   Check Gun Pulser
 Check Linac
  
  69   Mon 05 Mar 2012 00:03:33 Somkiat Wisaiphion Sun 04 Mar 2012 18:10:00 RepairLINAC Note Linac   Still no current at CM1 of Linac   
  68   Sun 04 Mar 2012 18:22:50 Somkiat Wisaiphion Sun 04 Mar 2012 18:00:00 RepairLINAC Note Linac  - Gun grid bias voltage OK 31.2 V by local
adjust to -31.2 V
  
  67   Sun 04 Mar 2012 18:20:46 Somkiat Wisaiphion Sun 04 Mar 2012 16:38:00 ProblemLINAC Note Linac  - Linac no current
- Check Linac (Gun grid bias voltage
error can not control by remote)
  
  66   Sun 04 Mar 2012 18:16:53 Somkiat Wisaiphion Sun 04 Mar 2012 16:36:00 RoutineOperation Note Warm Up Linac      
  65   Sun 04 Mar 2012 18:16:13 Somkiat Wisaiphion Sun 04 Mar 2012 16:17:00 RoutineOperation Note Warm Up Magnets   Warm up all magnets   
  64   Sun 04 Mar 2012 18:15:34 Somkiat Wisaiphion Sun 04 Mar 2012 15:42:00 ProblemMagnet Note Magnet  Reset water cooling system and magnets   
  63   Sun 04 Mar 2012 18:12:00 Somkiat Wisaiphion Sun 04 Mar 2012 15:38:00 ProblemElectrical Note Electrical 698- Electricity sag-swell (failure)
- PWP failure
- All water cooled magnet fault
  
  62   Sun 04 Mar 2012 08:34:21 Sudjakom Jummunt Sun 04 Mar 2012 08:20:00 RoutineOperation Note SPS Close Orbit correction    SPS close orbit correction
 2012-03-04_8-24.PNG2012-03-04_8-26.PNG 
  61   Sun 04 Mar 2012 08:32:29 Sudjakom Jummunt Sun 04 Mar 2012 08:15:00 RoutineOperation Note Ramping energy to 1.2Gev   Ramping Complete @ 131.1 mA  
Forward Power 14.4 kW
  
  60   Sun 04 Mar 2012 08:31:20 Sudjakom Jummunt Sun 04 Mar 2012 08:12:00 RoutineOperation Note Start Ramping      
  59   Sun 04 Mar 2012 08:31:00 Sudjakom Jummunt Sun 04 Mar 2012 08:11:00 RoutineOperation Note Stop Injection   Stop Injection @ 132.2 mA   
  58   Sun 04 Mar 2012 08:30:18 Sudjakom Jummunt Sun 04 Mar 2012 08:08:00 RoutineOperation Note Start Injection      
  57   Sun 04 Mar 2012 08:29:56 Sudjakom Jummunt Sun 04 Mar 2012 08:01:00 RoutineOperation Note Warm Up Magnets   Warm up HBT , STR magnets & Flush
TGP 1 cycle
  
  56   Sun 04 Mar 2012 08:28:45 Sudjakom Jummunt Sun 04 Mar 2012 08:00:00 RoutineOperation Note Dump Beam   Dump Beam @ 50.4 mA   
  55   Sat 03 Mar 2012 20:21:46 Ekwarong Amornwittawat Sat 03 Mar 2012 20:19:52 RoutineBeam Dynamics Note SPS Close Orbit correction    SPS close orbit correction
 2012-03-03_20-19.PNG 
  54   Sat 03 Mar 2012 20:16:16 Ekwarong Amornwittawat Sat 03 Mar 2012 20:14:12 RoutineOperation Note Ramping energy to 1.2Gev   Ramping Complete @132mA   
Accelerator Technology Division | Synchrotron Light Research Institute (Public Organization) | Powered by ELOG V3.1.1